OpenStudy (anonymous):

Determine if the class intervals are appropriate for the data below, or if they would result in too many classes. (Yes or No). 25, 32, 18, 99, 43, 16, 29, 35, 26, 24, 21, 33, 26, 17, 40, 22, 38, 16, 19. 1. 1 2. 10 3. 2

OpenStudy (anonymous):

start from 10...taking difference of 10 units...u'll get 10 classes...

OpenStudy (anonymous):

Sorry, but I don't know what you mean. Could you explain it all out?

OpenStudy (anonymous):

the least value in this distribution is 16...so start from 10...like this: class intervals: 10-20 20-30 30-40 40-50 50-60 60-70 70-80 80-90 90-100

OpenStudy (anonymous):

Thank you. Alright, then what?

OpenStudy (anonymous):

after that...in the interval 10-20...u have 16,16,18 and 19....so frequency is 4 like dat...complete the data for every interval...

OpenStudy (anonymous):

10-20 : 16, 16, 18, 19 20-30 : 21, 24, 25, 26 30-40 : 32, 33, 35, 38 40-50 : 40, 43, 60-70 : 70-80 : 80-90 : 90-100 : 99

OpenStudy (anonymous):

ya right...bt wat ws the qstn????

OpenStudy (anonymous):

Determine if the class intervals are appropriate for the data or if they would result in too many classes. Yes or no.

OpenStudy (anonymous):

and what are the options...the class interval should be 10...so if 10 is there..then that is the answer...

OpenStudy (anonymous):

Never mind. I read that wrong :P If 10 is there?

OpenStudy (anonymous):

Sorry, I'm a dummy when it comes to math

OpenStudy (anonymous):

bt this is statistics...is 10 in any of the options???

OpenStudy (anonymous):

1 10 2 Between these 3 I have to say yes or no.

OpenStudy (agent0smith):

I guess the 10 is the only one that works well. If you used 2 as the class interval for example, 16-18: 16, 16 18-20: 18, 19 20-22: 21 22-24: 24 24-26: ... and so on. Since your data ranges from 16 all the way up to 99, using a class interval 2 would give a lot of classes. You can use this: http://www.wyzant.com/Images/Help/Data_image_1.gif to see how many classes you'll have, but you don't need to - you can see from the table i wrote above that there'll be too many.

OpenStudy (anonymous):

So, for both 1 & 2, I would say yes... because they result in too many classes? The wording of the question is confusing

OpenStudy (agent0smith):

Yep, a class interval 1 and 2 will result in too many classes... you can use that formula above to see how many: Highest value - lowest value: 99-16 = 83 Divide that by the class interval 2: 83/2 = 41.5 classes (i guess that's too many...) Divide by the class interval 1: 83/1 = 83 classes (too many again)

OpenStudy (agent0smith):

Oh and yes, the question is confusingly worded... Determine if the class intervals are appropriate for the data below, or if they would result in too many classes. (Yes or No). Does this mean say "yes, the class interval is appropriate" or "Yes, too many classes"...